# Experimental Analysis of Efficiency of the Messaging Layer Security for Multiple Delivery Services

### v1

@@ Abstract
Messaging Layer security (MLS) and its underlying Continuous Group Key Agreement (CGKA) protocol allows a group of users to share a cryptographic secret in a dynamic manner, such that the secret is modified in member insertions and deletions. One of the most relevant contributions of MLS is its efficiency, as its communication cost scales logarithmically with the number of members. However, this claim has only been analysed in theoretical models and thus it is unclear how efficient MLS is in real-world scenarios. Furthermore, practical decisions such as the chosen Delivery Service and paradigm can also influence the efficiency and evolution of an MLS group. In this work we analyse MLS from an empirical viewpoint: we provide real-world measurements for metrics such as commit generation and processing times and message sizes under different conditions. In order to obtain these results we have developed a highly configurable environment for empirical evaluations of MLS through the simulation of MLS clients. Among other findings, our results show that computation costs scale linearly in practical scenarios even in the best-case scenario.
@@
Continuous Group Key Agreement , MLS , Delivery Service , Network simulation , Decentralised Communications
journal: Computers & Security
[…]
{highlights}
We study the efficiency of the Messaging Layer Security (MLS) protocol in an experimental setting.
We provide an implementation of a configurable environment for the empirical evaluation of MLS.
We measure latency, generation and processing times and message sizes for different Delivery Services, paradigms and group states.
@@ 1 Introduction
Messaging Layer Security (MLS) [1] is a recent communications standard for establishing secure messaging groups between a set of users. It is mainly composed by a cryptographic scheme named Continuous Group Key Agreement (CGKA), whose objective is to distribute a shared secret between members of the group. One of the most relevant characteristics of CGKA protocols is its flexibility in group composition: the scheme allows insertions and deletions of members while ensuring the security properties of Forward Secrecy (FS) and Post-Compromise Security (PCS).
@@
However, most contributions to the CGKA literature only address the theoretical aspects of these protocols. While some works perform empirical measurements of their CGKA variants [10, 11, 12], they are often limited to small groups and specific scenarios. This prevents the analysis of practical considerations that also affect the development of MLS groups. Specifically, the Authentication and Delivery services defined in the MLS RFC have received very little attention [13], and are usually instantiated as abstract functionalities. The latter is relevant for the execution of the CGKA protocol, as it handles the transmission of messages between members of a group and stores information required to insert new members. The treatment of these services as abstract and efficient obfuscate their influence on the correct functioning and efficiency of the protocol. Thus, the interaction between the underlying CGKA schemes and the real-world instantiation of these services has not been explored.
Furthermore, MLS introduces new elements to CGKA groups whose impact has not been studied in the literature. In particular, MLS defines a method for External Joins, through which users can enter a group without requiring invitation. The KeyPackage and GroupInfo messages are required to add new members to the group, and they need to be stored and made available to users. The propose-and-commit paradigm requires the generation, distribution and processing of proposal messages that represent an additional overhead.
In this work we address those issues by performing an empirical evaluation of the performance of MLS under multiple different scenarios. We study how performance - mainly commit generation and processing times, as well as message size - is affected by different conditions such as the chosen Delivery Service and paradigm.
To that end we have developed a testbed for experimental evaluations of MLS. Our implementation consists of a simulated MLS client that autonomously joins groups and periodically publishes messages and updates, as well as two different Delivery Services to distribute MLS messages between the clients. The testbed is highly configurable and allows for the deployment of an arbitrary number of simulated clients. We publish our implementation as open-source so that it is available for researchers to perform their own analysis or even create new scenarios.
@@
    A testbed for experimental evaluations of MLS. We provide a Rust implementation [^1] Available at https://github.com/SDABIS/mls_experimental_analysis of a simulated MLS client that can be configured in detail to perform various tests. When executed in parallel, these clients will create MLS groups, add or remove other clients and exchange application messages, according to their parameters. We also implement two different Delivery Services: a centralised MQTT server and a distributed P2P-based publish-subscribe network.
@@
    An empirical analysis of the performance of MLS. We study MLS behaviour under multiple conditions: the type of Delivery Service employed, the number and behaviour of users and whether or not some MLS functionalities like External Joins and the propose-and-commit paradigm are being employed. We measure the computational cost of generating and processing updates, the size of exchanged messages and the time required to deliver them to all members. Among other findings, our results show that the expected logarithmic complexity of MLS does not hold in practice, even in a best-case scenario.
The rest of the document is organised as follows: Section 2 will review related works to provide the context to our contribution. Section 3 will introduce to the reader the concepts of CGKA and MLS that are relevant for our analysis. In Section 4 we identify the variables we will measure and the parameters that will impact them. Section 5 introduces our implementation of the simulated MLS clients and the Delivery Services and in Section 4.2 we present our measurements of said environment. In Section 7 we discuss our most relevant findings and overall implications of our results. Finally, Section 8 will conclude this document and discuss future improvements.
@@ 2 Related Work
CGKA protocols have been thoroughly defined in the literature [14, 15, 16], exploring the security properties of Forward Security (FS) and Post-Compromise Security (PCS) [14]. Since one of the main objectives of CGKA protocols is to provide efficient scaling with the number of users [5], they usually employ binary trees to represent the state of the group [17, 2, 18]. As such, CGKA protocols usually aim for logarithmic complexity in adds and removals, although the PCS requirement unavoidably causes its efficiency to degrade under certain conditions [19]. The most popular instantiation of a CGKA protocol is the Messaging Layer Security (MLS), which has recently been standardised as RFC 9420 [1]. MLS employs TreeKEM [2] as its CGKA protocol, and uses the shared cryptographic secret to perform Secure Group Messaging between its members.
The efficiency of CGKA protocols has mostly been discussed in a theoretical framework, without presenting experimental results. The authors of [6] discuss the bandwidth required to distribute messages to all members of the CGKA group, and propose introducing a server to more efficiently provide each member with the information they require. A similiar approach is taken in [7], which also optimises the protocol for post-quantum algorithms for public key encryption. The impact of the tree’s state in efficiency is taken into account. Other works further develop this analysis to estimate the communication cost (i.e., the amount of messages required) of healing the group after a compromise. Communication cost is estimated both for a generic CGKA protocol [3, 4, 5] and for specific schemes [20, 21, 22]. The authors of [12] also analyse the communication cost of the TreeKEM protocol by focusing on the shape of the ratchet tree.
The authors of [23] propose a Delivery Service that employs Reliable Broadcast and consensus mechanisms to distribute proposals and commits, respectively. They later presented an implementation [24], but without experimental measurements. The CGKA variant introduced in [8] is particularly suited for its implementation using a Blockchain as a decentralised Delivery Service, although the work is mainly theoretical and no experimental environment is discussed. Conversely, the authors of [25] do implement a CGKA environment using a blockchain for IoT devices, but employ Asynchronous Ratchet Trees [17], an outdated version of TreeKEM and thus not compatible with MLS. In [11] an alternative CGKA protocol that does not employ binary trees is presented and an implementation is provided for a simplified execution environment.
Experimental analysis of CGKA implementations is limited to works that present an specific CGKA variant such as [10, 11, 26], although their experimental settings are limited: they only consider groups with few members - up to 128 - and do not model user behaviour. In both cases their measurements show how commit generation times increase linearly as the number of users grow.
@@ 3.1 Continuous Group Key Agreement
A Continuous Group Key Agreement (CGKA) protocol is a scheme that allows a set of users to establish a common secret. This shared value changes dynamically with every modification of the state of the group, whether it be insertions or eliminations of users. In addition to the shared secret, each member possesses some private information such as a signing key pair linked to an identity credential. Each state of the group is called an epoch; whenever a member issues a modification to the state of the group through a commit message, a new epoch is created with a different shared secret. Throughout this document, we will refer to the group member who generates a commit as committer.
CGKA protocols are designed to resist the leakage of the shared secret, whether it be because of a failure in the implementation or the compromise of a member, whose secrets are stolen [15]. As a result, the most relevant security properties a CGKA protocol aims to provide are Forward Secrecy (FS) and Post-Compromise Security (PCS) which state that if the shared secret of an epoch is compromised, it should not be possible to obtain any secret information from previous and following epochs.
A commit may add a new member to the group or remove an existing one. Additionally, members can modify their individual state if they believe it has been compromised. This also refreshes the group’s state and may modify some of the group’s contextual information and parameters. We will refer to this actions as Add, Remove and Update, respectively.
@@ 3.2 Messaging Layer Security (MLS)
For the correct functioning of an MLS group, two abstract services are defined [27]: the Authentication Service (AS) and the Delivery Service (DS). The former is tasked with generating the member’s credentials and assisting in validating the identity of other members. The AS is considered to be an external entity to the protocol: for example, it could be instantiated as the current Public Key Infrastructure, in which X.509 certificates containing a public key are signed by Certificate Authorities that are trusted by default.
Figure 1: Message flow in MLS through the Delivery Service.
On the other hand, the Delivery Service is tasked with distributing messages and storing relevant information about both groups and users. The flow of information in MLS is shown in Figure 1: Whenever group members send a message, it is published to the Delivery Service, which is tasked with forwarding it to the recipients (usually the other group members).
@@ 3.3 Communication in MLS
    Handshake: Messages that apply or propose changes to the group’s state. Usually handshake messages are generated by group members, with the exception of External Joins, in which non-members insert themselves into the group. By default, handshake messages are encrypted with the group’s shared secret. Handshake messages fall into the following categories:
    • (a)
      Proposal: does not apply a change directly to the group, but instead communicates the intention to do so.
    • (b)
      Commit: contains one or more proposals (not necessarily generated by the committer) and applies them to the group. It also contains the updated path secrets that the receivers will employ to update their Ratchet Tree.
@@
In addition to delivering protocol messages, the DS must also store some information generated by users and groups such that it can be accessed by any party on demand. We will refer to this abstract functionality of the DS as Directory. There are two types of objects that the Directory must store:
  • 1.
    Key Packages: they are generated by users and contain cryptographic information required to insert them into an MLS group, including their credential and capabilities as well as an Hybrid Public Key Encryption (HPKE) [28] public key. Key Packages are included in Add proposals and cannot be reused.
  • 2.
    Group Information (GroupInfo): they contain the a group’s Ratchet tree and other contextual information that is required for performing External Joins. It also contains an HPKE public key whose secret counterpart is derived from the group’s shared secret, and thus is known by all members. GroupInfo packages can be generated by any group member and represent its state for an specific epoch.
Since MLS messages are encrypted, the DS does not need to be a trusted party. However, we remark that a malicious or compromised DS still has the capability of tracking group compositions, reordering or deleting messages or even deny service to a party.
@@ 3.4 Group Updates
As mentioned, a group’s state is modified by commits which can either add or remove members or update their individual state. While the two latter operations are fairly simple, inserting new members into an MLS group requires additional steps. They can either be added through Invitation - if a current member inserts them into the group - or External Join - if they insert themselves [13]. These two methods are represented in Figure 2.
Invitation
@@
Figure 2(a) shows the message flow for inviting new users to the group. As a prerequisite, the new user Bob must have uploaded a Key Package created by him to the Delivery Service’s Directory. Then, the group member member Alice obtains it and generates an Add proposal that includes the Key Package. When the proposal is committed, Alice employs the HPKE public key contained in Bob’s Key Package to encrypt the group’s secrets and includes them in a Welcome message directed to Bob, along other contextual information required to participate in the group.
All other members of the group then process Alice’s commit. Bob’s Key Package also included a leaf node which is inserted into the group’s TreeKEM after validating his credentials.
External Join
@@
In contrast to Add proposals, External Joins are initiated by the joiner, who is not a member of the group. The procedure is shown in Figure 2(b): First, the joiner Bob accesses the group’s GroupInfo, which was uploaded to the Directory by the group member Alice. Then, he generates a value that will be used to derive the new epoch’s shared secret and encrypts it with the group’s HPKE public key (contained in the GroupInfo) such that it is accessible only to members of the group. Finally, Bob generates a commit with the Join proposal - a new proposal type defined by MLS -, which is validated and processed by all members of the group. We remark that an MLS group can specify if External Joins are allowed.
4 Empirical evaluation of MLS
In this Section we will analyse the most relevant properties of the execution environment in which we will measure the behaviour of MLS using different Delivery Services. We will start by defining the parameters that will configure our testbed. Then, we will introduce the variables that we will measure, whose values will be predictably affected by the parameter configuration.
4.1 Parameters
The most relevant metric for our experiments is group size, that is, the current number of members of the MLS group. It is claimed that MLS’ efficiency scales logarithmically with group size [6, 20], and all of our measurements are is significantly influenced by it. We will also analyse how the other parameters introduced in this Section interact with group size.
In addition to group size, we will measure the influence of the following parameters:
4.1.1 Delivery Service
The Delivery Service can be instantiated as a centralised server that manages the MLS groups and relays messages to its participants. The use of a server help to ensure consistency as it can ensure that all parties receive all messages and in the same order. However, the server also represents a single point of failure and may introduce bandwidth limitations, as all messages need to go through a single machine. On the other hand, a distributed network of devices that autonomously transmit and forward messages between them can also serve as a Delivery Service. In addition to the distribution of messages, the distributed network can also serve as Directory by storing KeyPackages and GroupInfo messages.
4.1.2 Paradigm
Users apply modifications to an MLS group by issuing a commit, which updates the shared secret and certain nodes in the Ratchet Tree. This can lead to inefficiencies if multiple commits are generated consecutively, as the updated secrets are needlessly overwritten. Alternatively, a commit can simultaneously apply multiple modifications through the use of proposals. We refer to this two different approaches as paradigms.
More specifically, we will distinguish between these paradigms:
@@ 4.1.2 Paradigm
The propose-and-commit paradigm reduces the number of modifications to the Ratchet Tree that are required to perform the same functionality. However, it also introduces an additional overhead: the generation, distribution and processing of proposal messages. In the propose-and-commit paradigm, a commit that applies $n$ proposals requires the exchange of $n+1$ messages: each of the $n$ proposal messages and the commit message itself. This contrasts with the commit paradigm, in which every handshake message applies exactly one modification. We unify these results under the following Average Update cost ($auc$) metric:
$auc=\frac{cc+\sum_{i=0}^{n}cp_{i}}{n}$
where $cc$ and $cp_{i}$ are the costs of the commit and its $i$th proposal, respectively.
@@
We will also consider External Joins in our discussion of paradigms. Although External Joins are not incompatible with any of the other two paradigms - as removes and updates can still be either commit or propose-and-commit - they do represent a different method for adding new members to the group that can be compared with them.
4.1.3 Ratchet Tree State
Although the TreeKEM protocol aims to achieve to scale logarithmically as the number of users grow, this is only the ideal case: when all the tree branches, included intermediate nodes, are populated. These branches can be blanked by some operations, such as invitations of removals under certain conditions [18]. In the worst case (i.e. whenever all intermediate nodes are blanked), the group is structured as a list, which implies that the scaling becomes linear [15].
The amount and location of blanked nodes depends mostly on which users perform the updates. A group reaches its worst possible state if all updates are issued by the same user - the group creator -, as only the nodes in its path to the root will be populated. In contrast, the best state is reached when all members issue an update, without considering removes. This ensures that all intermediate nodes are populated. This can be achieved by only allowing the last user to join to issue updates.
We will model these scenarios by restricting who is able to issue updates: the First user for the worst state and the Last for the best case. We will compare them to a normal scenario in which updates are performed by Random users.

### v2

# arXiv:2502.18303v2[cs.CR] 27 Feb 2026

@@ Abstract
Messaging Layer Security (MLS) and its underlying Continuous Group Key Agreement (CGKA) protocol allows a group of users to share a cryptographic secret in a dynamic manner, such that the secret is modified in member insertions and deletions. One of the most relevant contributions of MLS is its efficiency, as its communication cost scales logarithmically with the number of members. However, this claim has only been analysed in theoretical models and thus it is unclear how efficient MLS is in real-world scenarios. Furthermore, practical considerations such as the chosen paradigm and the evolution of the group can also influence the performance of an MLS group. In this work we analyse MLS from an empirical viewpoint: we provide real-world measurements for metrics such as commit generation and processing times and message sizes under different conditions. In order to obtain these results we have developed a highly configurable environment for empirical evaluations of MLS through the emulation of MLS clients. Among other findings, our results show that computation costs scale linearly in practical settings even in the best-case scenario.
@@
Continuous Group Key Agreement , MLS , Delivery Service , Performance Evaluation , MQTT
@@ 1 Introduction
Messaging Layer Security (MLS) [12] is a recent communications standard for establishing secure messaging groups between a set of users. It is mainly composed of a cryptographic scheme named Continuous Group Key Agreement (CGKA), whose objective is to distribute a shared secret between members of the group. One of the most relevant characteristics of CGKA protocols is its flexibility in group composition: the scheme allows insertions and deletions of members while ensuring the security properties of Forward Secrecy (FS) and Post-Compromise Security (PCS).
@@
However, most contributions to the CGKA literature only address the theoretical aspects of these protocols. Although some works perform empirical measurements of their CGKA variants [11, 34, 17], they are often limited to small groups and specific scenarios. Furthermore, MLS introduces new elements to CGKA groups whose impact has not been studied in the literature. In particular, MLS defines a method for External Joins, through which users can enter a group without requiring invitation. The KeyPackage and GroupInfo messages are required to add new members to the group, and they need to be stored and made available to users. The propose-and-commit paradigm requires the generation, distribution and processing of proposal messages that represent an additional overhead.
In this work we address those issues by performing an empirical evaluation of the performance of MLS under multiple different scenarios. We analyse how measurements such as commit generation, processing and distribution times or message size are affected by different experimental conditions: the size of the group, the chosen paradigm, or the evolution of the group state.
To that end we have developed a testbed for experimental evaluations of MLS. Our implementation consists of a emulated MLS client that autonomously joins groups and periodically publishes messages and updates, as well as two different Delivery Services to distribute MLS messages between the clients. The testbed is highly configurable and allows for the deployment of an arbitrary number of emulated clients. We publish our implementation as open-source so that it is available for researchers to perform their own analysis or even create new scenarios.
@@
    A testbed for experimental evaluations of MLS. We provide a Rust implementation [^1] Available at https://github.com/SDABIS/mls_experimental_analysis of a emulated MLS client that can be configured in detail to adapt their behaviour. When executed in parallel, these clients will create MLS groups, add or remove other clients and exchange application messages, according to their parameters.
@@
    An empirical analysis of the performance of MLS. We define various experiments in which the emulated clients are configured to recreate MLS groups with different configurations. The experiments performed study the impact of MLS functionalities like External Joins and the propose-and-commit, as well as the degradation of the group’s state by blanked intermediate nodes. We measure the computational cost of generating and processing updates, the size of exchanged messages and the time required to deliver them to all members. Among other findings, our results show the relevance of practical considerations such as the chosen paradigm and persistent storage. Furthermore, we identify that the expected logarithmic complexity of MLS does not hold in practice even in a best-case scenario due to the high cost of managing the group state, whose size increases linearly as the number of members grow.
The rest of the document is organised as follows: Section 2 will review related works to provide the context to our contribution. Section 3 will introduce to the reader the concepts of CGKA and MLS that are relevant for our analysis. In Section 4 we describe our testbed and the experiments to be performed. Section 5 introduces our implementation of the emulated MLS clients and the Delivery Service and in Section 6 we present our measurements of said environment. In Section 7 we discuss our most relevant findings and overall implications of our results. Finally, Section 8 will conclude this document and discuss future improvements.
@@ 2 Related Work
CGKA protocols have been thoroughly defined in the literature [5, 8, 6], exploring the security properties of Forward Security (FS) and Post-Compromise Security (PCS) [5, 15]. Since one of the main objectives of CGKA protocols is to provide efficient scaling with the number of users [16], they usually employ binary trees to represent the state of the group [20, 14, 25]. The most popular instantiation of a CGKA protocol is the Messaging Layer Security (MLS), which has recently been standardised as RFC 9420 [12].
The efficiency of CGKA protocols has mostly been discussed in a theoretical framework, without presenting experimental results. The authors of [7] discuss the bandwidth required to distribute messages to all members of the CGKA group, and propose introducing a server to more efficiently provide each member with the information they require. A similar approach is taken in [22], which also optimises the protocol for post-quantum algorithms for public key encryption. The impact of the tree’s state in efficiency is taken into account. Other works further develop this analysis to estimate the communication cost (i.e., the amount of messages required) of healing the group after a compromise. Communication cost is estimated both for a generic CGKA protocol [10, 9, 16] and for specific schemes [4, 19, 23]. The authors of [17] also analyse the communication cost of the TreeKEM protocol by focusing on the shape of the ratchet tree.
The authors of [29] propose a Delivery Service that employs Reliable Broadcast and consensus mechanisms to distribute proposals and commits, respectively. They later presented an implementation [30], but without experimental measurements. The CGKA variant introduced in [3] is particularly suited for its implementation using a Blockchain as a decentralised Delivery Service, although the work is mainly theoretical and no experimental environment is discussed. Conversely, the authors of [1] do implement a CGKA environment using a blockchain for IoT devices, but employ ART [20], an outdated version of TreeKEM and thus not compatible with MLS. In [34] an alternative CGKA protocol that does not employ binary trees is presented and an implementation is provided for a simplified execution environment.
Experimental analysis of CGKA implementations is limited to works that present an specific CGKA variant such as [11, 34, 33], although their experimental settings are limited: they only consider groups with few members —up to 128— and do not model user behaviour. In both cases their measurements show how commit generation times increase linearly as the number of users grow.
@@ 3.1 Continuous Group Key Agreement
A Continuous Group Key Agreement (CGKA) protocol is a scheme that allows a set of users to establish a common secret. This shared value changes dynamically with every modification of the state of the group, whether it be insertions or eliminations of users. In addition to the shared secret, each member possesses some private information such as a signing key pair linked to an identity credential. Each state of the group is called an epoch; whenever a member issues a modification to the state of the group through a commit message, a new epoch is created with a different shared secret. Throughout this document, we will refer to the group member who generates a commit as committer. CGKA protocols are designed to resist the leakage of the shared secret by providing both Forward Secrecy (FS) and Post-Compromise Security (PCS) across epochs.
A commit may add a new member to the group or remove an existing one. Additionally, members can modify their individual state if they believe it has been compromised. This also refreshes the group’s state and may modify some of the group’s contextual information and parameters.
@@ 3.2 Messaging Layer Security (MLS)
For the correct functioning of an MLS group, two abstract services are defined [13]: the Authentication Service (AS) and the Delivery Service (DS). The former is tasked with generating the member’s credentials and assisting in validating the identity of other members. The Delivery Service is tasked with distributing messages and storing relevant information about both groups and users.
@@ 3.3 Communication in MLS
    Handshake: Messages that apply or propose changes to the group’s state. Handshake messages can either be Proposals that communicate the intention to apply a change to the group but do not directly apply it; or Commits that contain one or more proposals and apply them to the group.
@@
In addition to delivering protocol messages, the DS must also store some information generated by users and groups such that it can be accessed by any party on demand. We will refer to this abstract functionality of the DS as Directory. The Directory stores KeyPackages, which are generated by users and contain cryptographic information required to insert them into an MLS group. They are included in Add proposals. Group Information (GroupInfo) packages are also stored in the Directory, and they contain the group’s Ratchet tree and other contextual information that is required for performing External Joins.
@@ 3.4 Group Updates
As mentioned, a group’s state is modified by commits which can either add or remove members or update their individual state. Although the two latter operations are fairly simple, inserting new members into an MLS group requires additional steps. They can either be added through Invitation —if a current member inserts them into the group— or External Join —if they insert themselves [21].
@@
Figure 1(a) shows the message flow for inviting new users to the group. As a prerequisite, the new user Bob must have uploaded a Key Package created by him to the Delivery Service’s Directory. Then, the group member member Alice obtains it and generates an Add proposal that includes the Key Package. When the proposal is committed, Alice creates a Welcome message directed to Bob, which contains the information needed to participate in the group. All other members of the group then process Alice’s commit.
@@
In contrast to Add proposals, External Joins are initiated by the joiner, who is not a member of the group. The procedure is shown in Figure 1(b): First, the joiner Bob accesses the group’s GroupInfo, which was uploaded to the Directory by the group member Alice. Then, he generates a value that will be used to derive the new epoch’s shared secret and includes it in a Join proposal which then committed by Bob and processed by all members of the group.
4 Testbed Design
In this Section we describe the scope of our experimental analysis of the performance of MLS. We start by introducing the experiments that study different aspects of the evolution of an MLS group. Then, we discuss the measurements we will extract from our each of the experiments.
Our testbed is composed of two elements. Firstly, a number of identical emulated MLS clients that independently participate in MLS groups by creating or joining them and issuing and processing updates. Secondly, a centralised Publish/Subscribe Delivery Service that forwards messages to group members. In this architecture, users can publish messages to different topics such that they are received by all users that are subscribed to said topic. Figure 2 shows the handling of topics in the testbed. Every group has a different topic to which all current members are subscribed. Every user is subscribed to a welcome topic through which it receives all Welcome messages directed to them. Additionally, all groups have their own topic that all members are subscribed to, and every handshake and application message of that group is published through it.
Figure 2: Publish/Subscribe architecture in our testbed.
As mentioned, our testbed aims to study the empirical performance of MLS under different situations. To that end, we have defined a series of experiments, that are characterised by the specific configuration that is shared by all emulated clients. In each of them we aim to study the evolution of the group under specific conditions. Table 1 shows a summary of the experiments to be performed, which will be detailed in this Section.
Experiment Batch Paradigm Committer Target users Measurements
Commit Group Evolution: Paradigm Commit Random 10000 Gen and Proc Time Commit, Welcome and KP Size Latency
2 Prop Propose-and-Commit Gen and Proc Time Commit Size
4 Prop
8 Prop
External Join External Join
First Group Evolution: Tree Degradation Commit First Gen and Proc Time (Deep) Commit Size
Last Last
100 Users Stable Group Size Random 100 Gen and Proc Time Commit Size
750 Users 750
5000 Users 5000
Table 1: Summary of the performed experiments.
4.1 Group Evolution
Our first batch of experiments explores how measurements are affected by the evolution of the MLS group as it grows in size. All experiments described in this section will consist on the emulation of an MLS group to which new members are constantly being added. The operation to perform and the user that will perform it will depend on the specific experiment.
4.1.1 Group Updates Paradigm
In MLS, the following paradigms can be employed to issue updates to the group, including adding new members:
@@ 4.1.1 Group Updates Paradigm
The propose-and-commit paradigm reduces the number of modifications to the Ratchet Tree that are required to perform the same functionality. However, it also introduces an additional overhead: the generation, distribution and processing of proposal messages.
@@
We will also consider External Joins in our discussion of paradigms. Although External Joins are not incompatible with any of the other two paradigms —as removes and updates can still be either commit or propose-and-commit— they do represent a different method for adding new members to the group that can be compared with them.
To study the impact of the chosen paradigm in MLS, we will conduct a total of five experiments that differ on how updates are issued and how new members are inserted into the group: one for the commit paradigm, one for External Joins and three for the propose-and-commit paradigm using different amounts of proposals per commit: 2, 4 and 8. We will refer to these experiments as Commit, External Join, 2 Prop, 4 Prop and 8 Prop, respectively.
4.1.2 Tree Degradation
(a) A member adds another user, which creates new empty nodes to the tree. As the committer cannot know the private keys of the new nodes, they remain empty.
(b) A member removes another user. To ensure the removed user cannot decrypt future updates, all of its known nodes are emptied.
Figure 3: Events that degrade the state of the Ratchet Tree.
The tree structure of the group state allows MLS to achieve logarithmic efficiency in the best-case scenario. However, this only occurs when all the intermediate nodes are populated. Figure 3 shows two scenarios in which intermediate nodes may become blanked: during additions to the rightmost side of the tree and removals. Blanked intermediate nodes are populated again only when a member whose path to the root includes it issues an update. The impact of blanked intermediate nodes is felt on the amount of ciphertext that need to be distributed, as shown in Figure 4.
(a) Ratchet tree on its optimal state, with all intermediate nodes populated. New commits includes 3 ciphertexts.
(b) Ratchet tree on its least optimal state, with all intermediate nodes empty. A new commit includes 7 ciphertexts.
Figure 4: Impact of tree degradation in the performance of MLS for a 8-member group.
We will analyse the importance of the tree degradation by conducting two different experiments. They will measure performance for the worst —all intermediate nodes empty — and best — all intermediate nodes populated— cases. The worst possible state is achieved if all updates are issued by the first user to join the group, as shown in Figure 3(a). Conversely, the best possible state is reached if every user issues an update; we achieve this by ensuring that only the last member to join can issue updates. We refer to this experiments as First (worst possible state) and Last (best possible state).
4.2 Stable Group Size
Whereas in our first batch of experiments we analyse how the measurements change as the group evolves in size, the second batch focuses on the performance of group operations at static group sizes. Our objective is to measure the variance of the cost of updates at a particular group size. Furthermore, we also analyse if this variance remains stable throughout different group sizes, or if larger groups have more volatile computing costs than smaller ones.
The experiments will unfold as follows: the group evolves normally until it reaches the target group size. Then, members will alternate between removing and adding users, such that the group size remains stable. We will perform three different experiments for target group sizes 100, 750 and 5000 users, as they represent different orders of magnitude close to a power of 2. We refer to this experiments as 100 Users, 750 Users and 5000 Users, respectively.